How many liters of a 30% solution should be added to 30 liters of a 10% solution if the final mixture is to contain a 24% solution?
Answer by mananth(16946)

........ Percent ----------------quantity
Solution A 10 ----------------30
Solution B 30 ----------------x
Total 24 ---------------- 30+x
10*30+30*x=24(30+x)
300+30x =720+24x
30x-24x =720-300
6x=420
/6
x=70 liters 30% Solution B